OpenStudy (anonymous):

In a polynomial division problem, the dividend does not have an x^3-term. When the problem is written in long division form, what should be the coefficient of the x^3-term in the dividend? A. 1 B. 0 C. -1 D. 3

OpenStudy (fibonaccichick666):

Do we get to see the problem?

OpenStudy (anonymous):

That is the problem @fibonaccichick666

OpenStudy (fibonaccichick666):

ok, so let's think about this then

OpenStudy (fibonaccichick666):

first, can you tell or show me what a dividend is?

OpenStudy (anonymous):

yes, its the certain amount that you want to divide up dividend/divisor=quotient

OpenStudy (fibonaccichick666):

ok, cool, now just one quick note before we continue, division is just fractions, so keep fraction rules and division rules in mind. However, this problem is much simpler than we think. Are you ok with me using a generic form for a polynomial or would you prefer a specific example?

OpenStudy (anonymous):

What ever works best.. Just be prepared to explain if i get confused. :P

OpenStudy (fibonaccichick666):

ok, why don't I do a specific example, if you are not in theoretical math, it will be much easier to understand.

OpenStudy (anonymous):

yes that works!!

OpenStudy (fibonaccichick666):

So let us try doing \(x^4+2x^2+3\) divided by \(x^3+x^2+1\)

OpenStudy (fibonaccichick666):

Can you just try setting it up for me?

OpenStudy (fibonaccichick666):

Try setting it up using our old long division format

OpenStudy (fibonaccichick666):

let me know what you get or if you are confused

OpenStudy (fibonaccichick666):

Just set it up, don't try to solve it

OpenStudy (anonymous):

ok sorry my internet went down.. so let me read what you said real quick

OpenStudy (fibonaccichick666):

it's cool, I just thought you were trying to solve it

OpenStudy (anonymous):

it would be \[\left( x^4+2x^2+3 \right) \div \left( x^3+x^2+1 \right)\]

OpenStudy (anonymous):

You mean like that?

OpenStudy (fibonaccichick666):

not quite, so long division style means (assume the square root is just a division thingy, I don't actuall know what it is called...) \[(x^3+x^2+1)\sqrt{x^4+2x^2+3}\]

OpenStudy (anonymous):

poop I was close.

OpenStudy (fibonaccichick666):



OpenStudy (fibonaccichick666):

you were, and this way of writing these is only used when doing poly div.

OpenStudy (fibonaccichick666):

so anyway, that is the first part

OpenStudy (fibonaccichick666):

so all this wants to know is how else we can write \(x4+2x2+3\) so that it "contains" an \(x^3\) term

OpenStudy (fibonaccichick666):

whoopsie \(x^4+2x^2+3\)

OpenStudy (fibonaccichick666):

so what does that mean? It means we sort of add a ghost term in, one that is there but the things are still equal with or without it.

OpenStudy (fibonaccichick666):

so we start by adding in the x^3 term, with an unknown coeficient

OpenStudy (fibonaccichick666):

\(x^4+ax^3+2x^2+3\)

OpenStudy (fibonaccichick666):

So now, we have hit the meat and potatoes of your question, what value of a makes it the exact same as \(x^4+2x^2+3\)

OpenStudy (anonymous):

1?

OpenStudy (fibonaccichick666):

What do you think a is?

OpenStudy (fibonaccichick666):

well, if a=1, we have \[x^4+ax^3+2x^2+3=x^4+1x^3+2x^2+3\] is that the same as \(x^4+2x^2+3\)?

OpenStudy (anonymous):

no?

OpenStudy (fibonaccichick666):

correct, no they are not the same

OpenStudy (fibonaccichick666):

the x^3 term will not go away

OpenStudy (fibonaccichick666):

we need an a value where the x^3 term goes away or vanishes

OpenStudy (anonymous):

i think we can eliminate 1 and -1 right?

OpenStudy (fibonaccichick666):

nope, because we only can put 1 variable in

OpenStudy (fibonaccichick666):

forget the multiple choice

OpenStudy (fibonaccichick666):

just be logical

OpenStudy (anonymous):

3?

OpenStudy (fibonaccichick666):

5+?=5 or x+?=x think about what the question mark would be

OpenStudy (anonymous):

0

OpenStudy (fibonaccichick666):

what does the x^3 term have to equal in order for the polynomial to remain the same

OpenStudy (fibonaccichick666):

you are guessing

OpenStudy (fibonaccichick666):

provide me with why

OpenStudy (anonymous):

Ok I think its 0 because. 5+?=5 if you take 5+3 = 8 it equals 8 as 5+0=5

OpenStudy (fibonaccichick666):

ok, so, yes the ? will =0. It is due to the additive identity if we want to be technical

OpenStudy (fibonaccichick666):

so how does that help us?

OpenStudy (fibonaccichick666):

any ideas?

OpenStudy (anonymous):

uhm we want something that'll will get rid or keep it the same?

OpenStudy (fibonaccichick666):

ok, so whatever we add must equal what?

OpenStudy (fibonaccichick666):

what can we add to any number and it will remain the same?

OpenStudy (anonymous):

x^3?

OpenStudy (fibonaccichick666):

no, we cannot add x^3 to any number and have it be the same

OpenStudy (anonymous):

oh oh oh 0

OpenStudy (anonymous):

brainfart moment

OpenStudy (fibonaccichick666):

ok, there we go(and it happens to all of us)

OpenStudy (fibonaccichick666):

so we need to add zero somehow, no

OpenStudy (fibonaccichick666):

we have not arrived at your answer yet

OpenStudy (anonymous):

oh xD ok

OpenStudy (fibonaccichick666):

now, we need to add zero, but we also need to add an x^3 term, so we need to find a value of a such that 0=ax^3

OpenStudy (fibonaccichick666):

Once you figure out what value of a works, you will have your answer. As I do not think you fully understand, please check this reference. An example is located about halfway down the page. http://www.purplemath.com/modules/polydiv3.htm